# Sculptor: Empowering LLMs with Cognitive Agency via Active Context Management

### v1

@@ 2 Related Work
Effectively processing long contexts remains a critical challenge for LLMs. Early efforts focused on expanding context windows through architectural improvements (Beltagy et al., 2020; Su et al., 2023; Chen et al., 2023). Subsequently, a substantial body of work sought to further optimize performance by augmenting LLMs with external memory systems, employing comprehensive memory architectures and multi-agent frameworks to overcome context limitations (Li et al., 2025c; Yang et al., 2024; Li et al., 2025b; Wang and Chen, 2025; Chhikara et al., 2025; Yu et al., 2025). The push for longer and more complex context processing led to the development of specialized evaluation benchmarks, such as NIAH (Kamradt, 2023),NeedleBench (Li et al., 2025a), RULER (Hsieh et al., 2024a),LongBench-v2 (Bai et al., 2025), MRCR (Vodrahalli et al., 2024), and PI-LLM (Wang and Sun, 2025). These benchmarks were instrumental in revealing that despite architectural and memory enhancements, modern LLMs still perform poorly on information-sparse tasks. Among these, work such as PI-LLM further identified a deeper reason for this phenomenon: proactive interference, where earlier information in the context disrupts the processing of later, more relevant content (Wang and Sun, 2025). These documented failures on information-sparse tasks, coupled with the diagnosis of proactive interference, provide a strong motivation for our approach of active context management. Unlike external memory solutions that focus on storage and retrieval, or traditional attention that uniformly processes all tokens, our method provides the model with explicit tools to selectively retain, compress, or ignore information directly within its working memory, thereby mitigating interference without altering the underlying architecture.
@@
The integration of external tools to augment LLM capabilities is a burgeoning field of research, designed to overcome inherent model limitations such as knowledge cutoffs, hallucination, and weak mathematical reasoning. Pioneering work in this area has largely followed two paradigms. On one hand, models like Toolformer (Schick et al., 2023) demonstrate that LLMs can be fine-tuned to learn when and how to call external APIs, seamlessly incorporating their outputs into the generation process. On the other hand, prompting-based frameworks like ReAct (Yao et al., 2023) show that LLMs can synergize chain-of-thought reasoning with tool use in a zero-shot manner, interleaving thought, action, and observation steps to solve complex tasks. Subsequent research has focused on improving the reliability and scope of tool use, with work like Gorilla (Patil et al., 2023) developing models specialized for accurate API invocation, and frameworks like ART (Paranjape et al., 2023) creating programmatic pipelines for tool-augmented multi-step reasoning. However, a common thread in this existing literature is the focus on using tools to interact with the external world—accessing calculators, search engines, or code interpreters. Sculptor diverges from this trend by proposing a novel class of tools for internal context management. Instead of augmenting the LLM with external knowledge, we empower it with cognitive tools to actively curate its own working memory. This positions our work as complementary to existing tool-use research. Our approach directly targets cognitive bottlenecks like proactive interference, rather than solely addressing knowledge or computational limitations.
3 Methodology
3.1 The Sculptor Framework for Active Context Management
3.2 The Sculptor Tool Suite
We equip LLMs with eight fundamental tools organized into four functional categories.
(2) Summary, Hide, and Restore involves five complementary tools for dynamic content management. summary_fragment generates focused AI-powered summaries of specific fragments using configurable LLM models, while revert_summary restores summarized content back to its original form, ensuring no information is permanently lost. fold_fragment hides fragment content while preserving its existence, displaying only a folded marker with character count to dramatically reduce visual clutter. expand_fragment reveals previously folded content when it becomes relevant again, enabling dynamic focus management throughout conversations. restore_context provides a complete reset mechanism that clears all fragment states and returns the conversation to its original form.
(3) Intelligent Search and Retrieval is accomplished through search_context, a unified interface supporting both exact matching and semantic search modes across user messages, assistant responses, or all content, with configurable similarity thresholds using OpenAI embeddings. get_search_detail retrieves extended context around search results, with the model specifying the desired surrounding character count. By appending search results to the end of conversation history, this approach mitigates the “lost in the middle” problem (Liu et al., 2023) where models struggle to locate information buried within long contexts.
3.3 Teaching LLMs to Use Sculptor Tools
We explore two approaches for enabling LLMs to effectively utilize the Sculptor tools. (1) Zero-shot tool calling leverages the inherent tool-calling capabilities of state-of-the-art models like Claude-4-Sonnet and GPT-4.1, which demonstrate strong zero-shot generalization abilities for function calling. These models can understand and execute our Sculptor tools without any specific training, relying on their pre-trained understanding of tool usage patterns and natural language descriptions of tool schema. To encourage consistent tool engagement, we set tool_choice=“required” for the first round of multi-turn conversations.
(2) Multi-turn RL training involves reinforcement learning where models learn optimal tool usage strategies through iterative feedback and reward signals. This approach aims to develop more sophisticated tool usage patterns and better timing decisions for when to apply different ACM operations. While this training approach shows promising potential for further performance improvements, it is currently a work in progress and results are not yet available for evaluation.
4 Experiments
Table 1: PI-LLM task results for selected models and their ACM tools variants.
Model Update Count Overall
2 4 8 16 32 64 128 256
Context Length (tokens)
0.5K 1K 2K 4K 8K 16K 32K 64K Average
Claude-4-Sonnet 99.87 99.13 95.65 92.17 84.78 81.74 65.22 69.57 86.02
Claude-4-Sonnet (w/ ACM Tools.) 79.57 90.43 91.74 98.26 92.17 91.74 87.39 77.83 88.64
GPT-4.1 100.00 96.96 91.30 79.57 67.83 63.04 63.91 50.43 76.63
GPT-4.1 (w/ ACM Tools.) 98.26 92.17 89.13 93.04 83.91 76.09 64.35 60.43 82.17
Deepseek-V3-0324 100.00 95.22 85.65 70.00 63.91 33.04 32.17 21.74 62.72
Deepseek-V3-0324 (w/ ACM Tools.) 53.48 73.91 90.00 79.13 37.39 53.04 55.65 11.74 56.79
We explore the effectiveness of Sculptor through direct tool calling without specific training. Currently, even with no targeted training, through appropriate prompt engineering and leveraging LLMs’ inherent tool calling generalization capabilities, we achieve improved performance on information-sparse tasks. Results from multi-turn RL training approaches are ongoing work and will be updated in future versions.
4.1 Evaluated Models and Benchmarks
We conduct preliminary evaluations on benchmarks that test proactive interference and long-context capabilities. (1) PI-LLM (Wang and Sun, 2025) tests proactive interference by continuously updating key-value pairs, measuring how well models can forget outdated information and focus on current mappings. We set update counts from 2 to 256 and average the results, with 46 update keys following the paper’s recommended settings, repeating each test 5 times. (2) NeedleBench (Li et al., 2025a) evaluates information retrieval and reasoning in varying information densities by requiring models to find and connect multiple pieces of information scattered throughout long documents. We select the Multi-Needle Reasoning task with fixed depth of 40, testing across context lengths of 1k, 2k, 16k, 64k, and 128k tokens, averaging results over 10 runs per dataset.
4.2 Evaluation Results
4.2.1 Main Results
Table 2: Multi-Needle Reasoning Task Results of NeedleBench-128K.
Model 2-Needle 3-Needle 4-Needle 5-Needle Overall
Claude-4-Sonnet 96.0 82.0 54.0 36.0 67.0
Claude-4-Sonnet (w/ ACM Tools.) 100.0 98.0 88.0 90.0 94.0
GPT-4.1 90.0 64.0 30.0 8.0 48.0
GPT-4.1 (w/ ACM Tools.) 96.0 84.0 60.0 44.0 71.0
DeepSeek-V3-0324 88.0 68.0 28.0 16.0 50.0
DeepSeek-V3-0324 (w/ ACM Tools.) 92.0 58.0 50.0 32.0 58.0
Tables 1 and 2 present the comprehensive evaluation results across PI-LLM and NeedleBench benchmarks. The results demonstrate overall improvements when models are equipped with Sculptor tools, particularly on tasks suffering from proactive interference or requiring multi-needle reasoning.
Figure 2: NeedleBench Multi-Needle Reasoning performance of Claude-4-Sonnet across different context lengths. The model with Sculptor tools (blue) consistently outperforms the vanilla model (purple), with performance gaps widening as needle count increases.
Figure 3: Average performance comparison of Claude-4-Sonnet across all needle counts. Sculptor tools provide substantial improvements, particularly at longer context lengths where the vanilla model struggles significantly.
4.2.2 Performance Analysis by Benchmark
PI-LLM Results: Sculptor tools provide improvements on the PI-LLM benchmark for most models. Claude-4-Sonnet and GPT-4.1 achieve gains of 2.62 and 5.54 points respectively, while DeepSeek-V3 shows a decrease of 5.93 points, indicating varying generalization capabilities to unseen tools across models. Tool usage occasionally leads to score degradation at specific update counts - Claude-4-Sonnet drops from 99.87 to 79.57 at update count 2, and DeepSeek-V3 shows an even larger drop from 100.00 to 53.48, as models sometimes over-fold useful information. Despite these variations, two out of three models show overall improvements, demonstrating successful transfer of tool-calling capabilities to long-context tasks. Models primarily leverage fragment_context and fold_fragment to “sculpt away” proactive interference.
NeedleBench Results: On multi-needle reasoning tasks, all Sculptor-enhanced models achieve improvements across different needle counts, as shown in Figures 2 and 3. Claude-4-Sonnet, GPT-4.1, and DeepSeek-V3 achieve gains of 27.0, 23.0, and 8.0 points respectively, with Claude-4-Sonnet reaching 90% accuracy even on 5-needle tasks. Unlike PI-LLM where models prefer fragment_context and fold_fragment, multi-needle reasoning tasks predominantly trigger search_context usage for rapid needle localization, explaining the consistent improvements across all models.
5 Limitations and Future Work
Although Sculptor achieves significant improvements on multiple benchmarks, it introduces a computational trade-off: active context management reshapes the input context, invalidating traditional prefix-based KV cache mechanisms and potentially increasing computational costs. We anticipate future infrastructure optimizations, including adaptive caching strategies, will mitigate this issue.
Additionally, current LLMs rely on generalization from tool-use training in other domains to operate ACM tools, which does not guarantee stable or correct usage. Future work includes leveraging reinforcement learning to train LLMs to effectively and selectively use active context management tools across diverse benchmarks, developing advanced tool scheduling algorithms to minimize cache invalidation, and exploring architectural changes to enable partial cache reuse.
Acknowledgments
We sincerely thank Chupei Wang and Jiaqiu Vince Sun for valuable discussions on proactive interference, and Yongting Zhang for discussions on tool design.

### v2

@@ Appendix A Related Work
Effectively processing long contexts remains a critical challenge for LLMs. Early efforts focused on expanding context windows through architectural improvements (Su et al., 2023; Chen et al., 2023; Beltagy et al., 2020) and sparse attention mechanisms (Yuan et al., 2025; Gao et al., 2025; Lu et al., 2025). Subsequently, a substantial body of work sought to further optimize performance by augmenting LLMs with external memory systems, employing comprehensive memory architectures and multi-agent frameworks to overcome context limitations (Li et al., 2025c; Yang et al., 2024; Li et al., 2025b; Wang & Chen, 2025; Chhikara et al., 2025; Yu et al., 2025). The push for longer and more complex context processing led to the development of specialized evaluation benchmarks, such as NIAH (Kamradt, 2023),NeedleBench (Li et al., 2025a), RULER (Hsieh et al., 2024a),LongBench-v2 (Bai et al., 2025), MRCR (Vodrahalli et al., 2024b), and PI-LLM (Wang & Sun, 2025). These benchmarks were instrumental in revealing that despite architectural and memory enhancements, modern LLMs still perform poorly on information-sparse tasks. Among these, work such as PI-LLM further identified a deeper reason for this phenomenon: proactive interference, where earlier information in the context disrupts the processing of later, more relevant content (Wang & Sun, 2025). These documented failures on information-sparse tasks, coupled with the diagnosis of proactive interference, provide a strong motivation for our approach of active context management. Unlike external memory solutions that focus on storage and retrieval, or sparse attention that modifies token processing patterns, our complementary method provides the model with explicit tools to selectively retain, compress, or ignore information directly within its working memory, thereby mitigating interference while operating alongside existing architectural enhancements.
@@
The integration of external tools to augment LLM capabilities is a burgeoning field of research, designed to overcome inherent model limitations such as knowledge cutoffs, hallucination, and weak mathematical reasoning. Pioneering work in this area has largely followed two paradigms. On one hand, models like Toolformer (Schick et al., 2023) demonstrate that LLMs can be fine-tuned to learn when and how to call external APIs, seamlessly incorporating their outputs into the generation process. On the other hand, prompting-based frameworks like ReAct (Yao et al., 2023b) show that LLMs can synergize chain-of-thought reasoning with tool use in a zero-shot manner, interleaving thought, action, and observation steps to solve complex tasks. Subsequent research has focused on improving the reliability and scope of tool use, with work like Gorilla (Patil et al., 2023) developing models specialized for accurate API invocation, and frameworks like ART (Paranjape et al., 2023) creating programmatic pipelines for tool-augmented multi-step reasoning. However, a common thread in this existing literature is the focus on using tools to interact with the external world—accessing calculators, search engines, or code interpreters. Sculptor diverges from this trend by proposing a novel class of tools for internal context management. Instead of augmenting the LLM with external knowledge, we empower it with cognitive tools to actively curate its own working memory. This positions our work as complementary to existing tool-use research. Our approach directly targets cognitive bottlenecks like proactive interference, rather than solely addressing knowledge or computational limitations.
From External Compression to Internal Context Curation
A complementary line of research focuses on reducing the computational and memory burden of long contexts through intelligent compression and selection mechanisms. The LLMLingua series (Jiang et al., 2023; 2024a; Pan et al., 2024) pioneered the use of smaller models as compressors, performing extractive compression to remove task-irrelevant sentences and phrases while preserving information density. LongLLMLingua (Jiang et al., 2024a) further advanced this approach with question-aware coarse-to-fine compression and dynamic compression ratios, achieving significant improvements on long-context benchmarks. Similarly, Selective Context (Li et al., 2023) formalizes context selection as a relevance-based filtering problem for reading comprehension tasks.At the inference level, several methods optimize KV cache management to handle longer sequences more efficiently. StreamingLLM (Xiao et al., 2024) introduces attention sink mechanisms for online processing of extremely long inputs, while Scissorhands (Liu et al., 2023b) selectively retains only the KV pairs that will be referenced in future computations. More recent work like SnapKV (Li et al., 2024) and KVQuant (Hooper et al., 2025) focus on pre-computation importance estimation and low-bit quantization respectively to achieve memory-efficient inference.While these compression and selection methods effectively reduce computational overhead, they share a fundamental limitation: the compression decisions are made externally to the reasoning process, either by separate models or fixed heuristics. This can lead to information loss that the primary LLM might deem crucial for its reasoning chain. In contrast, Sculptor enables the LLM itself to make context management decisions dynamically based on its internal reasoning state, ensuring that compression and selection align with the model’s cognitive needs rather than external approximations.
Revisable Generation and Editable Thought Processes
Autoregressive decoding makes early errors “sticky” causing later tokens to amplify rather than fix them. Proactive interference tests show that retrieval accuracy degrades as semantically related but obsolete updates accumulate, underscoring the cost of an immutable context (Wang & Sun, 2025). Causally, pruning failed reasoning branches—or removing their surface forms from the visible history—immediately improves subsequent correctness, indicating the harmful persistence of erroneous traces (Feng et al., 2025). To mitigate this prefix lock-in, one line of work explores parallel or branched reasoning, such as the search-based Tree of Thoughts and the native parallelism in ParaThinker, which reduces “tunnel vision” at a small latency overhead (Yao et al., 2023a; Wen et al., 2025). A more fundamental approach alters the generation process itself, making outputs inherently revisable. This includes models that perform discrete edits, like iterative refinement via masking (Ghazvininejad et al., 2019) or sequence modification through insertion and deletion operations (Stern et al., 2019; Gu et al., 2019). Another family of non-autoregressive paradigms, such as diffusion LMs, enables global backtracking by denoising entire sequences in parallel (Li et al., 2022; Austin et al., 2021). Complementing these architectural shifts, test-time self-revision loops like Self-Refine and Reflexion demonstrate that lightweight edits to intermediate outputs reliably improve final solutions (Madaan et al., 2023; Shinn et al., 2023). Collectively, these findings build a strong case for equipping models with mechanisms to remove, rewrite, or compress their working context during reasoning—rather than merely appending tokens—so they can correct course instead of being trapped by early errors.
Appendix B Baseline Model Performance
We evaluate M3 across standard benchmarks and compare it with both smaller-scale models (Qwen3-8B, Qwen3-14B) and frontier models to establish baseline capabilities before ACM training. The results are presented in Table 3. Despite being a 13B model, M3 demonstrates exceptional tool-use performance, achieving 61.0% on Tau2-retail (Barres et al., 2025) (vs. 27.9% for Qwen3-8B), 61.8% on AceBench (Chen et al., 2025) (vs. 24.3% for Qwen3-8B), and 32.0% on SWE-bench Verified (vs. 3.3% for Qwen3-8B). This strong tool-calling foundation makes it particularly suitable for demonstrating ACM effectiveness. Benchmark results for DeepSeek-V3, GPT-4.1, and Claude-4-Sonnet (excluding NeedleBench-MRS and PI-LLM) are taken from the Kimi-K2 technical report (Team et al., 2025).
For Qwen3-8B and Qwen3-14B models, we followed the official documentation [^1] https://qwen.readthedocs.io/en/latest/deployment/vllm.html to enable 128K context length support through RoPE scaling (Su et al., 2023) with the YaRN method (Peng et al., 2023), using a scaling factor of 4.0 to extend from their original 32K context window to 128K tokens. This configuration was necessary for fair comparison on long-context benchmarks.
Table 3: Performance of M3 (13B parameters) compared to other models on standard benchmarks. Left: smaller-scale models (8B-14B). Right: frontier models. M3 demonstrates particularly strong tool-use capabilities (Tau2, AceBench) and coding performance (SWE-bench Verified).
Benchmark Qwen3-8B Qwen3-14B M3 Kimi-K2 DeepSeek-V3 GPT-4.1 Claude-4-Sonnet
Coding Tasks
LiveCodeBench v6 (Pass@1) 50.2 51.8 25.1 53.7 46.9 44.7 48.5
MultiPL-E (Pass@1) 70.4 77.0 72.4 85.7 83.1 86.7 88.6
SWE-bench Verified (Pass@1) 3.3 5.8 32.0 51.8 36.6 40.8 50.2
Tool Use Tasks
Tau2 retail (Avg@4) 27.9 36.2 61.0 70.6 69.1 74.8 75.0
Tau2 airline (Avg@4) 18.0 39.0 54.0 56.5 39.0 54.5 55.5
AceBench (Acc.) 24.3 23.7 61.8 76.5 72.7 80.1 76.2
Math & STEM Tasks
MATH-500 (Acc.) 92.2 95.4 80.2 97.4 94.0 92.4 94.0
AIME 2024 (Avg@64) 60.9 60.8 17.7 69.6 59.4 46.5 43.4
GPQA-Diamond (Avg@8) 53.0 58.1 45.2 75.1 68.4 66.3 70.0
General Tasks
MMLU (EM) 80.1 85.0 78.6 89.5 89.4 90.4 91.5
MMLU-Pro (EM) 74.5 77.5 65.2 81.1 81.2 81.8 83.7
IFEval (Prompt Strict) 34.9 35.5 77.1 89.8 81.1 88.0 87.6
SimpleQA (Correct) 6.7 8.8 7.4 31.0 27.7 42.3 15.9
Appendix C Evaluation Details
C.1 Benchmark-Specific System Prompts
As described in Section 4.1, we employed prompt engineering to enhance tool utilization capabilities across frontier models. The system prompts presented here are the final optimized versions used in our zero-shot evaluation for PI-LLM and NeedleBench benchmarks. These benchmark-specific prompts significantly improved Claude-4-Sonnet’s performance, demonstrating how targeted prompt optimization can unlock more effective Sculptor tool usage patterns.
System Prompt for PI-LLM Benchmark System Prompt: You are an intelligent assistant specialized for PI-LLM (Proactive Interference) testing. Your task is to track continuous updates to multiple key-value pairs and accurately remember the latest value for each key amidst substantial interference information. Remember: First use the fragment_context tool to split the long text into multiple fragments, then use fold_fragment to fold unimportant, earlier key-value updates, allowing you to concentrate on the final updates. The recommended approach is to divide the entire update stream into multiple fragments (e.g., ten fragments), then keep only the last two or three fragments while folding the rest. This strategy enables focus on the current, most recent content without being distracted by earlier information.
Figure 7: System prompt for PI-LLM benchmark, designed to handle proactive interference through strategic tool usage.
System Prompt for NeedleBench Multi-Needle Reasoning System Prompt: You are an agent skilled at analyzing family relationships between different people. You have ”search_context” and ”get_search_detail” tools. You excel at conducting chained searches for key information in long texts until you find complete information to reach your desired final answer. When searching for the oldest ancestor, ensure that every person name found has been verified through the search tools to confirm they truly have no higher-level ancestors before concluding your reasoning.
Figure 8: System prompt for NeedleBench Multi-Needle Reasoning, optimized for multi-hop retrieval tasks.
C.2 Unified System Prompt
The unified system prompt is used in our initial zero-shot evaluations and RL training experiments. This minimal, general-purpose prompt provides only basic guidance about available capabilities without prescriptive task-specific strategies. As described in Section 4.1, experiments with this unified prompt revealed inherent challenges of unguided tool use, including suboptimal tool selection patterns and insufficient execution depth. During RL training, the same prompt enables the model to autonomously discover optimal tool usage patterns across diverse contexts, as shown in Figure 9.
Unified System Prompt System Prompt: You are a helpful assistant. You can autonomously manage your own context: fold irrelevant information, focus on useful details, summarize long texts to keep your context concise, and use search tools to find key information in large documents.
Figure 9: The unified general-purpose prompt used both in initial zero-shot evaluations (to understand natural tool interaction patterns) and in RL training (to enable autonomous learning of tool usage strategies without prescriptive guidance).
This approach ensures that the model learns generalizable context management strategies rather than memorizing task-specific patterns, leading to more robust performance across diverse long-context.
C.3 Benchmark Details
We provide detailed configurations for the benchmarks used in our experiments:
NeedleBench Multi-Needle Reasoning:
For efficiency while maintaining representativeness, we test with a fixed depth of 40%, as our tool-based approach shows minimal sensitivity to needle position within the context. We examine context lengths of 1k, 2k, 16k, 64k, and 128k tokens, with each configuration evaluated across 10 runs per dataset to ensure statistical significance. The multi-needle variant requires connecting 2, 3, 4, and 5 needles simultaneously, making it substantially more challenging than single-needle retrieval tasks.
Data Processing for Context Length Constraints:
To ensure evaluation within our model’s 128k context window, we apply minimal preprocessing. For MRCR, we filter out test samples exceeding 128k tokens. For LongBench v2, we truncate samples exceeding 128k tokens using our tokenizer.
C.4 Baseline Method Evaluation Details: RAG and MemAgent
We include two baseline methods for long-context processing in our main results (see Table 2): retrieval-augmented generation (RAG) and MemAgent. Both baselines are evaluated under a unified, lightweight interface that accepts plain strings or standard message arrays, without dataset-specific restructuring. To avoid introducing additional external capabilities, no external embedding models are used.
For RAG [^2] https://github.com/THUDM/LongBench/tree/main/LongBench/retrieval/BM25, we adopt a BM25-only pipeline aligned with LongBench-style retrieval. The input is sentence-split with the same punctuation and length heuristics as common LongBench implementations, then chunked at $200$ tokens. A pseudo query is formed by concatenating the first and last $500$ tokens of the full context when an explicit query is not provided. Chunks are ranked by BM25 and concatenated from high to low until the accumulated length reaches $\approx 1500$ tokens. The system prompt constrains the model to answer strictly based on the retrieved context. This BM25-only design avoids external dense embeddings and evaluates the model’s intrinsic ability to reason over the retrieved snippets.
For MemAgent(Yu et al., 2025) [^3] https://github.com/BytedTsinghua-SIA/MemAgent/blob/main/quickstart.py, we follow their implementation with iterative memory updates. Extremely long inputs are first symmetrically trimmed to a maximum visible length of about $120\,\text{k}$ tokens to avoid one-sided truncation. The remaining text is processed in fixed $5\,\text{k}$-token chunks. At each step the model updates an explicit ”memory” that preserves previously useful information and integrates newly relevant details from the current chunk; the final answer is generated using the last memory along with the query. When no explicit query is given, we construct a short pseudo query from the first and last $500$ tokens of the source. Unless otherwise noted, defaults are: max context length $\approx 120\,\text{k}$ tokens, chunk size $5\,\text{k}$ tokens, and maximum generation length $1024$ tokens.
These choices emphasize reproducibility and minimize confounds: BM25-only retrieval for RAG, and a standard memory-update routine for MemAgent, both avoiding external vector indices or proprietary interfaces. Detailed hyperparameters are reflected in the text above rather than bespoke configuration tables to keep the protocol concise and focused.
Appendix D Dynamic Context-Aware Training Data Collection
Algorithm 1 presents our conditional trajectory collection algorithm with incremental loss assignment for dynamic context-aware RL training. The algorithm identifies context-modifying tool calls and creates separate training instances at each modification point, with incremental loss assignment to prevent redundant learning across multiple trajectory snapshots.
Algorithm 1 Conditional Trajectory Collection with Incremental Loss Assignment
1:Initial query $Q$, complete interaction trajectory with assistant completions $\{C_{i}\}_{i=0}^{n}$ and tool results $\{T_{i}\}_{i=0}^{n-1}$.
2:Set of context-modifying tools $\mathcal{T}_{ctx}$ (fragment, fold, summarize, restore operations).
3:Training dataset $\mathcal{D}_{train}$ containing (trajectory, loss_mask) pairs.
4:Initialize $\mathcal{D}_{train}\leftarrow\emptyset$
5:Initialize $trained\_indices\leftarrow\emptyset$ $\triangleright$ Track completion indices that have been assigned loss=1
6:for $i=0$ to $n$ do
7:  Extract tool call $a_{i}$ from completion $C_{i}$
8:  if $a_{i}\in\mathcal{T}_{ctx}$ or $i=n$ then $\triangleright$ Context modification or final completion
9:   $\triangleright$ Create trajectory snapshot up to current point
10:   $trajectory\leftarrow[Q,C_{0},T_{0},C_{1},T_{1},\ldots,C_{i}]$
11:   if $i<n$ then
12:     $trajectory\leftarrow trajectory+[T_{i}]$ $\triangleright$ Include tool result if not final
13:   end if
14:   $\triangleright$ Create incremental loss mask
15:   Initialize loss_maskloss\_mask with zeros for all elements in trajectorytrajectory
16:   for $j=0$ to $i$ do
17:     if $j\notin trained\_indices$ then $\triangleright$ Only assign loss to new completions
18:      $loss\_mask[C_{j}]\leftarrow 1$ $\triangleright$ Enable loss for completion $C_{j}$
19:      $trained\_indices\leftarrow trained\_indices\cup\{j\}$
20:     end if
21:   end for
22:   $\mathcal{D}_{train}\leftarrow\mathcal{D}_{train}\cup\{(trajectory,loss\_mask)\}$
23:   if $a_{i}\in\mathcal{T}_{ctx}$ then $\triangleright$ Update query for next iteration if context modified
24:     $Q\leftarrow\text{ApplyToolEffect}(Q,a_{i},T_{i})$ $\triangleright$ Apply context modification
25:   end if
26:  end if
27:end for
28:return $\mathcal{D}_{train}$
Appendix E Training Configuration
To ensure reproducibility and facilitate future research building upon our work, we provide the detailed training hyperparameters and hardware configuration used for GSPO training in Table 4. These settings represent the optimal configuration determined through extensive experimentation for training Sculptor-M3-RL with dynamic context-aware capabilities.
Table 4: GSPO training configuration.
Training Hyperparameters Hardware & Parallelism
Learning rate $1\times 10^{-6}$ GPU type NVIDIA H800 (80GB)
Training iterations 200 Total GPUs 128 (64 train, 64 rollout)
Clip ratio (lower) 0.0003 Tensor parallel (TP) 1
Clip ratio (upper) 0.0004 Pipeline parallel (PP) 4
KL penalty ($\alpha$) 0.0 Context parallel (CP) 16
LM regularization 0.1 Data parallel (DP) 4
Optimizer AdamW Max sequence length 128k tokens
Reward Design.
Our reward function for GSPO training is defined as:
[…]
This design encourages correct task completion while penalizing excessive tool usage, overly long trajectories, and malformed outputs.
Appendix F Sculptor Tool Suite Schemas
We provide the complete JSON schemas for all six core Sculptor tools, detailing their parameters and usage specifications. When tools like fold_fragment or summarize_fragment modify context content, the original text is temporarily stored in memory to enable complete restoration via restore_fragment. This ensures no information is permanently lost during context management operations.
"type": "function",
"function": {
"name": "fragment_context",
"description": "Fragment conversation content between specified markers into manageable pieces. Useful for breaking down long text sections for detailed analysis.",
"parameters": {
"type": "object",
"properties": {
"start_marker": {
"type": "string",
"description": "Start marker text to identify the beginning of content to fragment"
"end_marker": {
"type": "string",
"description": "End marker text to identify the end of content to fragment"
"num_fragments": {
"type": "integer",
"default": 5,
"minimum": 1,
"maximum": 20,
"description": "Number of fragments to create (default: 5)"
"role": {
"type": "string",
"enum": ["user", "assistant", "all"],
"default": "user",
"description": "Which role’s messages to search in (default: user)"
"required": ["start_marker", "end_marker"],
"additionalProperties": false
Figure 10: JSON schema for fragment_context tool: Fragments conversation content between markers.
"type": "function",
"function": {
"name": "fold_fragment",
"description": "Fold (hide) a conversation fragment to reduce visible context length. The content is preserved and can be expanded later.",
"parameters": {
"type": "object",
"properties": {
"fragment_id": {
"type": "string",
"description": "ID of the fragment to fold (e.g., ’f1a2b3’)"
"required": ["fragment_id"],
"additionalProperties": false
Figure 11: JSON schema for fold_fragment tool: Hides fragments to reduce context.
"type": "function",
"function": {
"name": "restore_fragment",
"description": "Restore a fragment to its original content from ACM storage. Works for both summarized and folded fragments.",
"parameters": {
"type": "object",
"properties": {
"fragment_id": {
"type": "string",
"description": "ID of the fragment to restore (e.g., ’f1a2b3’)"
"required": ["fragment_id"],
"additionalProperties": false
Figure 12: JSON schema for restore_fragment tool: Restores modified fragments.
"type": "function",
"function": {
"name": "summarize_fragment",
"description": "Summarize a conversation fragment using LLM to compress content while preserving key information. Supports focus-oriented summarization.",
"parameters": {
"type": "object",
"properties": {
"fragment_id": {
"type": "string",
"description": "ID of the fragment to summarize (e.g., ’f1a2b3’)"
"focus": {
"type": "string",
"description": "Focus area for the summary (e.g., ’technical details’, ’key decisions’, ’action items’, ’main points’, ’problems’, ’solutions’)"
"required": ["fragment_id", "focus"],
"additionalProperties": false
Figure 13: JSON schema for summarize_fragment tool: Compresses fragments with LLM.
"type": "function",
"function": {
"name": "search_context",
"description": "Search tool for finding exact text matches in conversation history.",
"parameters": {
"type": "object",
"properties": {
"query": {
"type": "string",
"description": "Exact text to search for in conversation history"
"role": {
"type": "string",
"enum": ["user", "assistant", "all"],
"default": "user",
"description": "Filter by message role (default: user)"
"max_results": {
"type": "integer",
"default": 10,
"minimum": 1,
"maximum": 50,
"description": "Maximum number of results to return"
"context_size": {
"type": "integer",
"default": 200,
"minimum": 50,
"maximum": 1000,
"description": "Context characters before/after match"
"required": ["query"],
"additionalProperties": false
Figure 14: JSON schema for search_context tool: Exact text search in conversation.
"type": "function",
"function": {
"name": "get_search_detail",
"description": "Get detailed context for a search result by its ID. Retrieves extended context around the search match position.",
"parameters": {
"type": "object",
"properties": {
"search_id": {
"type": "string",
"description": "Search result ID from search_context (e.g., ’s1a2b3’)"
"extended_context": {
"type": "integer",
"default": 500,
"minimum": 100,
"maximum": 2000,
"description": "Number of characters to show before and after the match (default: 500)"
"required": ["search_id"],
"additionalProperties": false
Figure 15: JSON schema for get_search_detail tool: Retrieves extended context.